OpenStudy (anonymous):

Horney little sloot looking for love in all the wrong places! 0=sqrt(x^2+ln(x)) solve for x then help me with this injective proof sqrt(x^2+ln(x))=sqrt(y^2+ln(y))

OpenStudy (anonymous):

answer the question and ill send you something special ;)

zepdrix (zepdrix):

Lol you're a creepy lil lady huh :p Hmm sorry my brain is too fried today to figure this one out :c

OpenStudy (anonymous):

i am what i am. so here's what i have for this.. really simple actually For every integers x and y, f(x)=f(y)---> x=y so, sqrt(x^2+ln(x))=sqrt(y^2+ln(y)) *end proof* The problem with the first part is that when trying to solve for x with wolfram i get the attached pic. Problem is is that I am not familiar with complex-valued logs

OpenStudy (anonymous):

i found the proof here http://courses.engr.illinois.edu/cs173/sp2009/lectures/lect_15_supp.pdf

zepdrix (zepdrix):

Oh that weird Product-Log thing? I've never really learned about that either.. hmm

OpenStudy (anonymous):

i guess ill read up on it then

zepdrix (zepdrix):

ya, soz :c

OpenStudy (aum):

Are you allowed to use a graphing calculator to solve 0=sqrt(x^2+ln(x)) ?

OpenStudy (anonymous):

yeah i don't know what to enter though

OpenStudy (aum):

0=sqrt(x^2+ln(x)) implies x^2+ln(x) = 0 You can plot x^2+ln(x) and find where it crosses the x-axis and that will be your solution OR you can plot y = x^2 and y = -ln(x) and see where they intersect.

OpenStudy (anonymous):

err nvm about the graphing party. anyway to do this numerically?

OpenStudy (aum):

Like Newton-Raphson method?

OpenStudy (anonymous):

ive never used it

OpenStudy (anonymous):

does the method use complex valued logs?

OpenStudy (aum):

Which chapter of the lesson does this question appear under. That may give an idea what method they are expecting you to use to solve this.

OpenStudy (aum):

No, Newton-Raphson method uses calculus. You start with a guess and then you successively approximate a better and better root until you reach the required level of accuracy.

OpenStudy (anonymous):

it's somewhere in the transcendental function section..

OpenStudy (aum):

But they would have discussed various methods for solving transcendental equations. If you list them we can choose the appropriate one to apply here.

OpenStudy (aum):

You can use a graphing method or use a numerical method. If they use numerical method we need to know what are allowed methods. But if you want to solve this by trial and error and a calculator I can tell you how.

OpenStudy (anonymous):

of ive done the ive just read about as newtons method

OpenStudy (kirbykirby):

I think Newton's Method would work nicely here. It will give you an approximate solution

OpenStudy (anonymous):

sweet

OpenStudy (aum):

I suppose some books call it Newton's Method and some others Newton-Raphson method. Have you covered calculus? Do you know how to find derivatives? Or do you want the trial and error method?

OpenStudy (aum):

Or do you want the trial and error method without calculus.

OpenStudy (anonymous):

yeah i've done it before in calc I but its a little new using natural logs. use calc please

OpenStudy (aum):

Guess two values for x where x^2+ln(x) changes sign (positive to negative or negative to positive). We are trying to solve x^2+ln(x) = 0 or x^2 = -ln(x) The LHS is always positive. Therefore, ln(x) has to be negative for the above equation to work. When is ln(x) negative? ln(x) is zero when x = 1. When x > 1, ln(x) is positive. Therefore, x has to be less than 1 for ln(x) to be negative. And we know ln(x) is not defined for x <=0. Therefore, out x must be 0 < x < 1. Pick an x value in the interval (0,1). Evaluate f(x). Pick another number so that you get a change in sign of f(x). Let me know when you get thiose two x values.

OpenStudy (aum):

Using your calculator, find f(0.5), f(0.25), f(0.75).

OpenStudy (aum):

where f(x) = x^2 + ln(x)

OpenStudy (anonymous):

.25=-1.324 .5=-.4431 .75=.27482 maybe .6?

OpenStudy (aum):

Therefore, the root is between x = 0.5 and 0.75 Try the middle value x = (0.5 + 0.75) / 2 = 0.625 (BTW, this trial and error method is called a binary search)

OpenStudy (anonymous):

ok i get the point so calc would be maybe .6-(x(x^2+ln(x)))

OpenStudy (anonymous):

then i just plug in what i got to the x subscript ns

OpenStudy (anonymous):

forever

OpenStudy (aum):

Binary search using the middle point often leads to the solution faster. Evaluate f(0.625) Does it change sign between x = 0.5 and 0.625 or between 0.625 and 0.75? Then you find the midpoint of that interval and keep going until the desired level of accuracy is reached.

OpenStudy (anonymous):

alright, thanks

OpenStudy (aum):

let us say the desired level of accuracy is two decimal places. Then you keep refining x until two successive solution has the same first two decimals and that is when you stop the iteration.

OpenStudy (aum):

you are welcome.